A group of 120 students, 7 have black hair, 35 have blond hair, 12 have red hair, and the remaining have brown hair. If a student is selected uniformly at random from the group, the probability that the student has black hair or blond hair can be expressed as a fraction a/ba/b, what is such fraction?
Answer #1

Let,

A denote the student have black hair

B denote a student have blond hair

C denote a student have red hair

D denote a student have brown hair

P[A]=7/120 = 0.0583

P[B] = 35/120 = 0.29167

P[C] = 12/120 = 0.1

P[D] = (120-7-35-12)/120 = 0.55

If a student is randomly selected then probability that the student has black hair or blond hair